OpenStudy (anonymous):

Help and explain ? 3√x + 3 =15

OpenStudy (anonymous):

3sqrt(x) + 3 = 15 3sqrt(x) = 12 sqrt(x) = 4 x = 16

OpenStudy (anonymous):

The idea is to get your variables (x) alone on one side, while the constants (15, 3, numbers) are on the other side. Using this idea we moved the 3 over to the other side. Divided 3 from both sides, and to get rid of the sqrt sign we squared both sides.

OpenStudy (anonymous):

Thankyou! but what if it's like this

OpenStudy (anonymous):

√6-3x-2=0

OpenStudy (anonymous):

the 6 and the 3x are both under the radical sign

OpenStudy (anonymous):

so the sqrt function is the same as saying everything under the radical sign to the 1/2 in otherwards, (6-3x)^1/2. To get rid of the 1/2 from under the radicand we can simply square both sides of the function. You see: sqrt(6-3x) - 2 = 0 sqrt(6-3x) = 2 6-3x = 4 (Here we squared both sides since you have to apply the operation to both sides for it to remain true) then we solve for x. -3x = -2 x = 2/3

OpenStudy (anonymous):

this helped alot, thanks for explaining!